OpenStudy (anonymous):

a • b • 4 = 4 • a • b is an example of which algebraic property? Associative Property of Multiplication Transitive Property Reflexive Property Commutative Property of Multiplication

OpenStudy (anonymous):

commutative property

OpenStudy (anonymous):

3 + x = 4 is the same as 4 = 3 + x. This is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

this is symmetric a = b means b = a

OpenStudy (anonymous):

symmetric: a=b implies b=a commutative: a+b = b+a associative: a+(b+c) = (a+b)+c

OpenStudy (anonymous):

transitive is if a=b and b=c then a=c